one apple and five oranges cost twenty nine rupees while three apples and four oranges cost thirty two rupees. find the total cost of four apples and five oranges.
Answer by mananth(16946)

Apple ----------x
Orenge ------------ y
1 x + 5 y = 29 .............1
3 x + 4 y = 32 .............2
Eliminate y
multiply (1)by -4
Multiply (2) by 5
-4 x -20 y = -116
15 x 20 y = 160
Add the two equations
11 x = 44
/ 11
x = Rs 4.00 Apple
plug value of x in (1)
1 x + 5 y = 29
4 + 5 y = 29
5 y = 29 -4
5 y = 25
y = Rs 5.00 Orange